# Physics-Enhanced Deep Surrogate for the Phonon Boltzmann Transport Equation

### v1

@@ 4.2 Accurate and Efficient Design
Since currently only the single-mode BTE is differentiable on OpenBTE, direct gradient-based optimization is not readily available for our problem. As a result, we include a non-gradient-based Bayesian Optimization (BO) [Jones1998EGO, Forrester2009RecentSurrogate] strategy as a strong and widely-used baseline for optimization of expensive-to-evaluate black-box functions. The BO routine proceeds by modeling the objective as a Gaussian Process over the binary domain and selecting points that balance exploration and exploitation using an acquisition function. Our acquisition function used is the Expected Improvement (EI) [Keane2006StatImprovement].
@@
Assuming the GP posterior at point $\mathbf{x}$ has predictive mean $\mu(\mathbf{x})$ and standard deviation $\sigma(\mathbf{x})$, and letting $f^{*}$ be the best (error minimizing) observed value so far, the EI can be computed in closed form as $\mathrm{EI}(\mathbf{x})=\left[\max(f^{*}-f(\mathbf{x}),0)\right]=(f^{*}-\mu(\mathbf{x}))\cdot\Phi\!\left(\tfrac{f^{*}-\mu(\mathbf{x})}{\sigma(\mathbf{x})}\right)+\sigma(\mathbf{x})\cdot\phi\!\left(\tfrac{f^{*}-\mu(\mathbf{x})}{\sigma(\mathbf{x})}\right)$, where $\Phi$ and $\phi$ denote the cumulative distribution function (CDF) and probability density function (PDF) of the standard normal distribution, respectively. An initial Sobol (quasi-random) phase can be introduced to fill up the space, providing a warm start for the optimization encouraging exploration [snoek2012practical]. We implemented the BO baseline using the high-level interface provided by the ax library [ax_platform] and set as convergence criterion a tolerance of 5% on the conductivity value. In Table 2 we present results for 8 example target thermal conductivities. These are non-uniformly spaced, reflecting the inherent skewness of the underlying distribution of the $30$ million candidate geometries. PEDS+AL achieves the best average design fractional errors at $4.0\%$, dominated by the material fabrication error. The other surrogates do not perform as well with PEDS alone at an average of $8.4\%$ and GP and GP+AL at $7.4\%$ and $9.0\%$, respectively. The current best (gradient-free) optimization uses the BTE solver and reaches the average of $2.4\%$. Plain PEDS exhibits substantially higher errors than PEDS+AL, underlining the role of active learning in improving surrogate robustness in more difficult regions of the space. Notably, the total design error that is the sum of model error and optimization error is dominated by the model error. In terms of computational efficiency, the surrogates are orders of magnitude faster than BTE. More specifically, for a single optimization run with PEDS evaluation takes $0.22$ seconds (almost 3 orders of magnitude faster). However, the strong batch-independence is so that for multiple parallel optimization runs (say 128 in this case) the costs is 0.002 (leading to 4 to 5 orders of magnitudes of improvement, 1 to 2 coming from the parallelism). Each design can be obtained within seconds to a couple of minutes, compared to several hours required for the full BTE optimization runs. Importantly, once the surrogate models are trained, they can be re-used for multiple design tasks at no additional training cost and at an overall cost comparable to solving one single BTE, amortizing the total training costs in only four optimization runs.
Kappa
Target
Model
Evaluations
Computation
Time
Kappa
Optimized
Relative
Model
Error
Design
Fractional
Error
@@ 5 Model Interpretability
PEDS offers a significant interpretability advantage over purely data-driven models, as the internal parameters of its low-fidelity solver inputs and outputs have clear physical meaning, enabling a direct assessment of the learning. A principal component analysis (PCA) of the predicted conductivities for the test geometries shows that the structure–transport relationship is intrinsically low-dimensional: the first two principal components capture 95% of the variance, revealing a clear trend governed by the effective conductivity (Fig. 5a). Two dominant modes emerge: one along which conductivity remains nearly constant, and another where it increases systematically. High-conductivity geometries occupy a wider region of the PCA space, reflecting greater structural diversity, whereas low-conductivity designs collapse into a tighter and more homogeneous cluster. To further interpret the model’s learned physics, we examine the relationship between the ground-truth BTE conductivity and the learned mixing coefficient $w_{\phi}$, normalized by the generator output norm $w_{\phi}\|G\|$ (Fig. 5b). This parameter controls how much the network corrects the low-fidelity Fourier geometry. Its nonlinear dependence on $\kappa_{BTE}$ demonstrates that the model has successfully learned the transition between transport regimes: large values of $w_{\phi}\|G\|$ correspond to low-conductivity (ballistic-dominated) cases where the Fourier solver alone is insufficient, while smaller coefficients characterize high-conductivity (diffusive) cases adequately captured by the linear model. The structure of this mapping mirrors the nonlinear patterns previously observed between $\kappa_{BTE}$ and $\kappa_{FE}$, confirming that the learned coefficient encodes a physically meaningful transition. In this sense, we can say that PEDS discovers the domain of validity of Fourier. Finally, representative generated geometries corresponding to low, intermediate, and high conductivities show only subtle visual differences, indicating that the latent space is smooth and effectively low-dimensional (Fig. 5c). The generator converges toward consistent pore configurations for each regime, supporting the hypothesis that the mapping between geometry and effective conductivity is governed by a few dominant structural modes.
@@
Physically, the Knudsen number (Kn) for a given geometry characterizes the transition between ballistic and diffusive transport regimes. This number effectively captures information about broadband MFP distributions in a complex geometry. Typically defined for single–MFP materials, we employ a recent computational method adapted for mode-resolved systems [hosseini2023reduced]. For each structure, the following equality holds:
@@
The Knudsen number measures the ratio of the mean free path (MFP) to a characteristic geometric feature length, distinguishing structures where the MFP exceeds the feature size ($\text{Kn}>1$, ballistic regime) from those where it is smaller ($\text{Kn}<1$, diffusive regime). In Fig. 6, the Knudsen number is computed for a set of representative geometries and shown to correlate strongly with the generated ballistic correction $w_{\phi}\|G\|$.
Figure 6: The generated ballistic correction $w\|G\|$ is highly linearly correlated with the Knudsen number of the geometry (dashed line). The higher the Knudsen number, the bigger is the MFP compared to the representative feature size and the more important will be the ballistic correction to the diffusive model. This confirms that PEDS is able to recover the transition between diffusive and ballistic transport and compensate for the ballistic transport.
@@ 6 Concluding remarks
Our results solidify the hypothesis, supported also by previous work [Pestourie_Mroueh_Rackauckas_Das_Johnson], that including a simple low‐fidelity solver can impart valuable inductive bias that substantially simplifies the learning task (Fig. 4). By removing the burden of enforcing the governing equations from the neural network, the incorporated low-fidelity physics allows it to focus only on correcting non‐diffusive phonon‐size‐effect errors. This explains its data efficiency: it inherits the low-data advantages of a physically informed baseline while avoiding the large data requirements typical of nerual operator for highly mode-resolved BTE problems [kennedy2000predicting, forrester2007multi, lu2019deeponet, li2020fourier]. Restricting the neural network’s role to geometry transformation aligns with our view that deep parametric models excel at learning features, representations, and low-dimensional projections, while numerical solvers are best suited for capturing the underlying physics. This stands in contrast to approaches that attempt to learn full solution operators, even when only a specific property of the solution is of interest [lu2022multifidelity].
@@
In fact, to compute the thermal conductivity, also accounting for its multi-fidelity version, the DeepONet in [lu2022multifidelity] required a significantly larger number of evaluations (10x with respect to our base experiment, almost 30x if we consider our active learning). The errors are not totally comparable as they are computed with respect to the $\kappa$ for PEDS and with respect to the solution field for the neural operator. Similarly to Ref. [lu2022multifidelity], PEDS can be seen as a natural extension of multi‑fidelity methods, hard‑coding a cheap physics core and learning only the input correction. To our knowledge, this is the first multifidelity scientific machine learning approach that achieves accuracy for BTE using a Fourier Equation low-fidelity solver. By learning the linear combination weight parameter between macro temperature field and a nano-scale residual, PEDS inherently bridges scales, making it intuitively well suited for problems where different approaches are used for different scales.
@@
Future work will focus on enriching the parameterization, making it bigger and continuous, and consequently, leverage a bigger neural network. In this work, we decided to keep the same parameterization as in [lu2022multifidelity]. This relatively small number of design parameters may also help explain the competitiveness of Gaussian processes. For both GPs and neural networks, scaling to richer continuous parameterization requires exponentially more data. This phenomenon, often referred to as the curse of dimensionality, is ubiquitous in machine learning and more broadly in the computational sciences. As the number of training points increases, the computational cost of fitting standard GPs becomes prohibitive, making classical GPs computationally expensive [rasmussen2006gaussian, wang2016gaussian] and likely less attractive as the results seem to show our results. In addition, their fitting cost increases as $O(N^{3})$ where $N$ is the number of training instances [lookman2019active], while neural network training cost grows linearly with N [rasmussen2006gaussian].
@@
Another practical motivation for a surrogate approach is to avoid repeated inverse design across different materials and scattering physics: Recent ML studies in phonon scattering and thermal-conductivity demonstrate that transfer learning across scattering-order approximations or material families can substantially reduce the number of expensive first-principles or BTE solves required to reach predictive accuracy [guo2023fast]; incorporating material descriptors in the learning process and applying transfer learning techniques [huisman2021survey, vettoruzzo2024advances] could allow our surrogates to extend their utility and to adapt to semiconductors other than silicon with relatively few additional high-fidelity solves [guo2023fast]. Another promising direction arises from operating in the input-space machine learning regime. Combining transferability strategies with a learned low-dimensional representation or a well-chosen set of geometric parameters can help mitigate the curse of dimensionality, enabling faster adaptation to new tasks [marzban2025hilab].
@@
We are currently exploring the use of the surrogate to accelerate the high-fidelity BTE solver itself, where PEDS or a neural operator can provide warm starts or learned preconditioners for iterative BTE solvers, similar to Ref. [kopanivcakova2025deeponet]. Applying these ideas to OpenBTE solvers could convert some of the cost of offline surrogate into online solver speedups and improve inverse-design throughput [li2023learningpreconditioners]. Beyond acceleration, the idea of preconditioning through learned surrogates naturally connects to transferability: a surrogate trained on one family of geometries or materials could initialize solvers for related configurations, accelerating convergence even when the underlying physics changes moderately.
@@
Although our approach was applied to the phonon BTE, the underlying methodology can be readily extended to the electron BTE [ziman2001electrons], where the the low-fidelity model is the drift-diffusion model [Chen2005].Additionally, this methodology may also extend to multiscale transport problems governed by physics analogous to the BTE. Examples include neutron transport [sun2024discrete] and rarefied gas dynamics [zhang2005lattice], when the fluid is diluted and the continuum Navier-Stokes cannot be applied. Exploring these other applications could unveil how these surrogate models encode scale-bridging inductive biases that generalize beyond phonon transport, providing a unified framework for a broader class of problems.

### v2

@@ 4.2 Accurate and Efficient Design
Since currently only the single-mode BTE is differentiable on OpenBTE, direct gradient-based optimization is not readily available for our problem. As a result, we include a non-gradient-based Bayesian Optimization (BO) [37, 23] strategy as a strong and widely-used baseline for optimization of expensive-to-evaluate black-box functions. The BO routine proceeds by modeling the objective as a Gaussian Process over the binary domain and selecting points that balance exploration and exploitation using an acquisition function. Our acquisition function used is the Expected Improvement (EI) [40].
@@
Assuming the GP posterior at point $\mathbf{x}$ has predictive mean $\mu(\mathbf{x})$ and standard deviation $\sigma(\mathbf{x})$, and letting $f^{*}$ be the best (error minimizing) observed value so far, the EI can be computed in closed form as $\mathrm{EI}(\mathbf{x})=\left[\max(f^{*}-f(\mathbf{x}),0)\right]=(f^{*}-\mu(\mathbf{x}))\cdot\Phi\!\left(\tfrac{f^{*}-\mu(\mathbf{x})}{\sigma(\mathbf{x})}\right)+\sigma(\mathbf{x})\cdot\phi\!\left(\tfrac{f^{*}-\mu(\mathbf{x})}{\sigma(\mathbf{x})}\right)$, where $\Phi$ and $\phi$ denote the cumulative distribution function (CDF) and probability density function (PDF) of the standard normal distribution, respectively. An initial Sobol (quasi-random) phase can be introduced to fill up the space, providing a warm start for the optimization encouraging exploration [85]. We implemented the Bayesian optimization (BO) baseline using the high-level interface of the ax library [65], minimizing the $l_{2}$ distance to the target conductivity, $\lvert\kappa(\mathbf{x})-\hat{\kappa}\rvert^{2}$, with convergence defined as reaching within 5% fractional error of the target value. In Table 2 we present results for 8 example target thermal conductivities. These are non-uniformly spaced, reflecting the inherent skewness of the underlying distribution of the $30$ million candidate geometries. The figures in the table are the model error (i.e. the fractional error between the conductivity predicted by the surrogate and the actual value for that geometry) and the design error (the discrepancy, in fractional error, between the desired thermal conductivity and value for the selected geometry). The design error is the sum of model error and optimization error. PEDS+AL achieves the best average design fractional errors at $4.0\%$, dominated by the material fabrication error. The other surrogates do not perform as well with PEDS alone at an average of $8.4\%$ and GP and GP+AL at $7.4\%$ and $9.0\%$, respectively. The current best (gradient-free) optimization uses the BTE solver and reaches the average of $2.4\%$. Plain PEDS exhibits substantially higher errors than PEDS+AL, underlining the role of active learning in improving surrogate robustness in more difficult regions of the space. Notably, the total design error that is the sum of model error and optimization error is dominated by the model error. In terms of computational efficiency, the surrogates are orders of magnitude faster than BTE. More specifically, for a single optimization run with PEDS evaluation takes $0.22$ seconds (almost 3 orders of magnitude faster). However, the strong batch-independence is so that for multiple parallel optimization runs (say 128 in this case) the costs is 0.002 (leading to 4 to 5 orders of magnitudes of improvement, 1 to 2 coming from the parallelism). Each design can be obtained within seconds to a couple of minutes, compared to several hours required for the full BTE optimization runs. Importantly, once the surrogate models are trained, they can be re-used for multiple design tasks at no additional training cost and at an overall cost comparable to solving one single BTE, amortizing the total training costs in only four optimization runs.
Kappa Target Model Evaluations Computation Time Kappa Optimized Relative Model Error Design Fractional Error
@@ 5 Model Interpretability
PEDS offers a significant interpretability advantage over purely data-driven models, as the internal parameters of its low-fidelity solver inputs and outputs have clear physical meaning, enabling a direct assessment of the learning. A principal component analysis (PCA) of the 25-dimensional mode-resolved conductivity fields (5Œ5 grids) predicted by the neural network reveals that the structure–transport relationship is intrinsically low-dimensional: the first two principal components capture 95% of the variance, revealing a clear trend governed by the effective conductivity (Fig. 5a). Two dominant modes emerge: one along which conductivity remains nearly constant, and another where it increases systematically. High-conductivity geometries occupy a wider region of the PCA space, reflecting greater structural diversity, whereas low-conductivity designs collapse into a tighter and more homogeneous cluster. To further interpret the model’s learned physics, we examine the relationship between the ground-truth BTE conductivity and the learned mixing coefficient $w_{\phi}$, normalized by the generator output norm $w_{\phi}\|G\|$ (Fig. 5b). This parameter controls how much the network corrects the low-fidelity Fourier geometry. Its nonlinear dependence on $\kappa_{BTE}$ demonstrates that the model has successfully learned the transition between transport regimes: large values of $w_{\phi}\|G\|$ correspond to low-conductivity (ballistic-dominated) cases where the Fourier solver alone is insufficient, while smaller coefficients characterize high-conductivity (diffusive) cases adequately captured by the linear model. The structure of this mapping mirrors the nonlinear patterns previously observed between $\kappa_{BTE}$ and $\kappa_{FE}$, confirming that the learned coefficient encodes a physically meaningful transition. In this sense, we can say that PEDS discovers the domain of validity of Fourier. Finally, representative generated geometries corresponding to low, intermediate, and high conductivities show only subtle visual differences, indicating that the latent space is smooth and effectively low-dimensional (Fig. 5c). The generator converges toward consistent pore configurations for each regime, supporting the hypothesis that the mapping between geometry and effective conductivity is governed by a few dominant structural modes.
@@
Physically, the Knudsen number (Kn) for a given geometry characterizes the transition between ballistic and diffusive transport regimes. This number effectively captures information about broadband MFP distributions in a complex geometry. Typically defined for single–MFP materials, we employ a recent computational method adapted for mode-resolved systems [33]. For each structure, the following equality holds:
@@
The Knudsen number measures the ratio of the mean free path (MFP) to a characteristic geometric feature length, distinguishing structures where the MFP exceeds the feature size ($\text{Kn}>1$, ballistic regime) from those where it is smaller ($\text{Kn}<1$, diffusive regime). In Fig. 6a, the Knudsen number is computed for a set of representative geometries and shown to correlate strongly with the generated ballistic correction $w_{\phi}\|G\|$. Although PEDS does not explicitly enforce similarity between coase temperature fields, it is indirectly able to reconstruct the coarse temperature fields from BTE. Using a representative set of geometries, we compare the temperature field obtained by solving the Fourier equation with the neural-corrected $5\times 5$ thermal conductivity tensor generated by PEDS against the temperature field obtained by downsampling the corresponding BTE solution. The median percentage error is $16.89\%$, which is primarily influenced by errors along the central line, where the true temperature approaches zero and small absolute deviations result in large relative errors. To provide a more robust assessment, we also report the mean absolute error, which is $0.053\,\mathrm{K}$. When normalized by the overall temperature range (approximately $1\,\mathrm{K}$), this corresponds to a relative error of $0.05$. These results show further generalization cpabilities, highlighting the ability of PEDS to recover key interpretable features of BTE thermal transport at the solution level in addition to the effective property.
Figure 6: (a) The generated ballistic correction $w_{\phi}\|G\|$ is highly linearly correlated with the Knudsen number of the geometry (dashed line). The higher the Knudsen number, the bigger is the MFP compared to the representative feature size and the more important will be the ballistic correction to the diffusive model. This confirms that PEDS is able to recover the transition between diffusive and ballistic transport and compensate for the ballistic transport. (b) Qualitative comparison of the downsampled BTE temperature and the one computed internally by PEDS for the pore geometry used in Fig.  2. This corresponds to an $8\%$ relative percentage error with respect to the BTE.
@@ 6 Concluding remarks
Our results solidify the hypothesis, supported also by previous work [71], that including a simple low‐fidelity solver can impart valuable inductive bias that substantially simplifies the learning task (Fig. 4). By removing the burden of enforcing the governing equations from the neural network, the incorporated low-fidelity physics allows it to focus only on correcting non‐diffusive phonon‐size‐effect errors. This explains its data efficiency: it inherits the low-data advantages of a physically informed baseline while avoiding the large data requirements typical of neural operator for highly mode-resolved BTE problems [41, 22, 59, 55]. Restricting the neural network’s role to geometry transformation aligns with our view that deep parametric models excel at learning features, representations, and low-dimensional projections, while numerical solvers are best suited for capturing the underlying physics. This stands in contrast to approaches that attempt to learn full solution operators, even when only a specific property of the solution is of interest [58].
@@
Similarly to Ref. [58], PEDS can be seen as a natural extension of multi‑fidelity methods, hard‑coding a cheap physics core and learning only the input correction. To our knowledge, this is the first multifidelity scientific machine learning approach that achieves accuracy for BTE using a Fourier Equation low-fidelity solver. By learning the linear combination weight parameter between macro temperature field and a nano-scale residual, PEDS inherently bridges scales, making it intuitively well suited for problems where different approaches are used for different scales.
@@
Future work will focus on enriching the parameterization, making it bigger and continuous, and consequently, leverage a bigger neural network. In this work, we decided to keep the same parameterization as in [58]. This relatively small number of design parameters may also help explain the competitiveness of Gaussian processes. For both GPs and neural networks, scaling to richer continuous parameterization requires exponentially more data. This phenomenon, often referred to as the curse of dimensionality, is ubiquitous in machine learning and more broadly in the computational sciences. As the number of training points increases, the computational cost of fitting standard GPs becomes prohibitive, making classical GPs computationally expensive [76, 92] and likely less attractive as the results seem to show our results. In addition, their fitting cost increases as $O(N^{3})$ where NN is the number of training instances [57], while neural network training cost grows linearly with N [76].
@@
Another practical motivation for a surrogate approach is to avoid repeated inverse design across different materials and scattering physics: Recent ML studies in phonon scattering and thermal-conductivity demonstrate that transfer learning across scattering-order approximations or material families can substantially reduce the number of expensive first-principles or BTE solves required to reach predictive accuracy [28]; incorporating material descriptors in the learning process and applying transfer learning techniques [35, 90] could allow our surrogates to extend their utility and to adapt to semiconductors other than silicon with relatively few additional high-fidelity solves [28]. Another promising direction arises from operating in the input-space machine learning regime. Combining transferability strategies with a learned low-dimensional representation or a well-chosen set of geometric parameters can help mitigate the curse of dimensionality, enabling faster adaptation to new tasks [63].
@@
We are currently exploring the use of the surrogate to accelerate the high-fidelity BTE solver itself, where PEDS or a neural operator can provide warm starts or learned preconditioners for iterative BTE solvers, similar to Ref. [46]. Applying these ideas to OpenBTE solvers could convert some of the cost of offline surrogate into online solver speedups and improve inverse-design throughput [54]. Beyond acceleration, the idea of preconditioning through learned surrogates naturally connects to transferability: a surrogate trained on one family of geometries or materials could initialize solvers for related configurations, accelerating convergence even when the underlying physics changes moderately.
@@
Although our approach was applied to the phonon BTE, the underlying methodology can be readily extended to the electron BTE [98], where the the low-fidelity model is the drift-diffusion model [15]. Additionally, this methodology may also extend to multiscale transport problems governed by physics analogous to the BTE. Examples include neutron transport [88] and rarefied gas dynamics [95], when the fluid is diluted and the continuum Navier-Stokes cannot be applied. Exploring these other applications could unveil how these surrogate models encode scale-bridging inductive biases that generalize beyond phonon transport, providing a unified framework for a broader class of problems.